You have chosen A. There are 3 possibilities for the prize, A, B, and C; we assume they are equally likely
If the prize is behind A, the emcee reveals B, the left-most of the 2 remaining doors. (p = 1/3)
If the prize is behind B, the emcee reveals C. (p = 1/3)
If the prize is behind C, the emcee reveals B. (p = 1/3_
Thus, the emcee reveals B in 2 cases, the prize is behind A and the prize is behind C. The total probability is 1/3+1/3=2/3
Thus, the probability of the prize being behind A if the emcee opens B = 1/3/(2/3) = 1/2, and the probability of the prize being behind C if the emcee opens B = 1/3/(2/3) = 1/2
The prize is equally likely to be behind A or C, so there is no advantage to making either choice.
Tom K.
06/25/16